# Stacked Pseudo-Convergent Sequences and Polynomial Dedekind Domains

### v2

@@ 2.1 Residually algebraic torsion extensions of $\mathbb{Z}_{p}$
Suppose now that $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$.
@@
Suppose now that $n\geq 2$ and assume by induction that $(s_{n-2},s_{n-1})$ is distinguished. Let $c\in\overline{\mathbb{Q}_{p}}$ be such that $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$. Since $(s_{n-2},s_{n-1})$ is distinguished, we have $v_{p}(s_{n-1}-c)\leq v_{p}(s_{n-1}-s_{n-2})=\delta_{n-2}<\delta_{n-1}$. Hence,
@@
We now show that $E^{\prime}=\{s_{n}\}_{n\in\mathbb{N}}$ is equivalent to $E=\{t_{n}\}_{n\in\mathbb{N}}$. Let $\{\lambda_{n}\}_{n\in\mathbb{N}}$ and $\{\delta_{n}\}_{n\in\mathbb{N}}$ be the gauges of $E$ and $E^{\prime}$, respectively. We need to show that, for each $k\in\mathbb{N}$, there exists $n\in\mathbb{N}$ such that $\lambda_{k}\leq\delta_{n}$. Since $E^{\prime}$ is unbounded, there exists $n\in\mathbb{N}$ such that $[\mathbb{Q}_{p}(t_{k}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. Since $(s_{n},\delta_{n})$ is a minimal pair, we have $v_{p}(s_{n}-t_{k})<\delta_{n}$, so that
@@
Finally, we need to show that, if $k\in\mathbb{N}$, then there exist $n_{0},m_{0}\in\mathbb{N}$ such that for each $n\geq n_{0}$ and $m\geq m_{0}$, we have $v_{p}(t_{n}-s_{m})>\lambda_{k}$. Let $n_{0}$ be the smallest integer such that $[\mathbb{Q}_{p}(t_{k}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n_{0}}):\mathbb{Q}_{p}]$. As in (2.6) above, $\lambda_{k}<v_{E}(X-s_{n_{0}})=\delta_{n_{0}}$. Let now $m>k$ and $n\geq n_{0}$. Then,
@@
Since $E$ is of transcendental type, for each $f\in\mathbb{Q}_{p}[X]$, $v_{E}(f)=v(f(s_{n}))$ for all $n$ sufficiently large (see §1.1). It follows that for each $\phi\in\mathbb{Q}_{p}(X)$, $\phi=f/g$, for some $f,g\in\mathbb{Q}_{p}[X]$, $v_{E}(\phi)=v_{E}(f)-v_{E}(g)$ is in $\Gamma_{n}$ for all $n$ sufficiently large. Hence, $\Gamma_{w}\subseteq\bigcup_{n}\Gamma_{n}$. Conversely, let $n\in\mathbb{N}$ and $f\in\mathbb{Q}_{p}[X]$ be of degree smaller than $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. Then, each root $\alpha_{i}$ of $f(X)$ in $\overline{\mathbb{Q}_{p}}$ has degree smaller than $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ and so, since $(s_{n},\delta_{n})$ is a minimal pair, we have
@@
Let now $n\in\mathbb{N}$ and $\overline{c}=\overline{f(s_{n})}\in k_{n}$, for some $f(s_{n})\in O_{n}^{*}$ where $f\in\mathbb{Q}_{p}[X]$ has degree strictly smaller than $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. In particular, $\overline{c}\not=0$. As in (2.10), $v_{E}(f(X))=v(f(s_{m}))=0$ for each $m\geq n$. Let $\alpha_{i}$ be a root of $f(X)$ in $\overline{\mathbb{Q}_{p}}$. Then by (2.9), $\overline{v}_{E}(X-\alpha_{i})=v_{p}(s_{n}-\alpha_{i})=v_{p}(d_{i})$ for some $d_{i}\in\overline{\mathbb{Q}_{p}}$. Then
@@
Conversely, let $\phi=f/g\in W\subset\mathbb{Q}_{p}(X)$, for some $f,g\in\mathbb{Q}_{p}[X]$. Let $\alpha_{i},\beta_{j}$ be the roots in $\overline{\mathbb{Q}_{p}}$ of $f$ and $g$, respectively. There exists $n\in\mathbb{N}$ such that $[\mathbb{Q}_{p}(\alpha_{i}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ and $[\mathbb{Q}_{p}(\beta_{j}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ for all $i$ and $j$. Hence, as in (2.9) we have
@@
    $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]$ and $\delta_{n}=\delta_{n}^{\prime}$.
@@
Without loss of generality, we may assume that $\overline{W}=\overline{\mathbb{Z}_{p}}_{E}=\overline{\mathbb{Z}_{p}}_{E^{\prime}}$.
Let $w$ be a valuation associated to $\overline{W}$.
We prove i). We have $s_{0},t_{0}\in\mathbb{Q}_{p}$. There exists $n\in\mathbb{N}$, $n\geq 1$, such that $w(X-s_{n-1})\leq w(X-t_{0})<w(X-s_{n})$, otherwise $t_{0}$ would be a pseudo-limit of $E$, which is not possible. In particular, $v_{p}(s_{n}-t_{0})=w(s_{n}-X+X-t_{0})=w(X-t_{0})\geq w(X-s_{n-1})=\delta_{n-1}$. If $n>1$ we have $[\mathbb{Q}_{p}(t_{0}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$ so by iii) of Definition 2.1 we have that $v_{p}(s_{n}-t_{0})=v_{p}(s_{n-1}-t_{0})<v_{p}(s_{n}-s_{n-1})=\delta_{n-1}$ which is impossible. Hence, $n=1$, so $v_{p}(s_{1}-t_{0})=w(X-t_{0})\geq w(X-s_{0})$. Reversing the roles of $s_{0},t_{0}$, we get the other inequality, so $w(X-s_{0})=w(X-t_{0})=\delta_{0}=\delta_{0}^{\prime}$.
Let $n\in\mathbb{N}$ and suppose that for each $m\leq n$ we have $[\mathbb{Q}_{p}(s_{m}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{m}):\mathbb{Q}_{p}]$ and $\delta_{m}=\delta_{m}^{\prime}$.
Since $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$, by ii) of Definition 2.1 we have $v_{p}(t_{n+1}-s_{n})\leq v_{p}(t_{n+1}-t_{n})=\delta_{n}^{\prime}=\delta_{n}$. Now,
@@
Now, if $[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$, then since $(s_{n},t_{n+1})$ is distinguished, we would have $v_{p}(s_{n+1}-t_{n+1})\leq v_{p}(t_{n+1}-s_{n})$, which is impossible. Hence, $[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]\geq[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$. The other inequality is proved in a symmetrical way, so $[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$.
@@
Let $n\geq 1$. If $f\in\mathbb{Q}_{p}[X]$ has degree strictly smaller than $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]$, then by (2.10) we have $w(f(X))=v_{p}(f(s_{n}))$ and also $w(f(X))=v_{p}(f(t_{n}))$, so $v_{p}(f(s_{n}))=v_{p}(f(t_{n}))$. This proves that $\Gamma_{n}=\Gamma_{n}^{\prime}$ and so $e_{n}=e_{n}^{\prime}$.
Suppose now that $f\in\mathbb{Q}_{p}[X]$ of degree strictly smaller than $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]$ is such that $v_{p}(f(s_{n}))=v_{p}(f(t_{n}))=0$. In particular, $w(f(X))=0$ by (2.10). By (2.11) and the analogous equation where $s_{n}$ is replaced by $t_{n}$, we get that $f(s_{n}),f(t_{n})$ have the same residue as $f(X)$, so in particular, $k_{n}=k_{n}^{\prime}$. Therefore, $f_{n}=f_{n}^{\prime}$. ∎
@@ 2.2 Residually algebraic extensions of $\mathbb{Z}_{p}$ which are DVRs
In this section we characterize DVRs of $\mathbb{Q}_{p}(X)$ extending $\mathbb{Z}_{p}$ such that the residue field extension is algebraic, necessarily of infinite degree by Corollary 2.12; this fact has already been noted in a different way in [21, p. 4217]. We will see in §2.3 that there is no such restriction on the residue field degree for DVRs of $\mathbb{Q}(X)$ which are residually algebraic extensions of $\mathbb{Z}_{(p)}$ (see Corollary 2.28).
@@
for $\text{for }\alpha\in\mathbb{C}_{p},\text{ we have }e_{\alpha}\cdot f_{\alpha}<\infty\text{ if and only if }\alpha\in\overline{\mathbb{Q}_{p}}.$
@@
If $K_{1}$ is a tame extension of $\mathbb{Q}_{p}$, then the ramification index of $L$ over $\mathbb{Q}_{p}$ is equal to lcm$\{e(K_{1}|\mathbb{Q}_{p}),e(K_{2}|\mathbb{Q}_{p})\}$ (see for example [7]), so $e$ divides $e_{1}$ and the claim is true.
@@
The following result is analogous to [21, Theorem 2.5].
@@
We show in Theorem 2.21 that given any algebraic extension $k$ of $\mathbb{F}_{p}$ and group $\Gamma$ such that $\mathbb{Z}\subseteq\Gamma\subseteq\mathbb{Q}$, there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $\mathbb{Z}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$, provided that either $[k:\mathbb{F}_{p}]$ is infinite or $\Gamma$ is not discrete (this condition being necessary by Corollary 2.12).
@@
Let $k$ be an algebraic extension of $\mathbb{F}_{p}$ and $\Gamma$ a totally ordered group with $\mathbb{Z}\subseteq\Gamma\subseteq\mathbb{Q}$, such that either $[k:\mathbb{F}_{p}]$ or $[\Gamma:\mathbb{Z}]$ is infinite (the last condition is equivalent to $\Gamma$ being not discrete). Then there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $k_{\alpha}=k$ and $\Gamma_{\alpha}=\Gamma$. In particular, $\mathbb{Z}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$.
Note that, by Corollary 2.12, the last claim shows that $[k:\mathbb{F}_{p}]\cdot[\Gamma:\mathbb{Z}]=\infty$ is necessary.
@@
Since $\overline{\mathbb{F}_{p}}$ is countable, we may suppose that $k=\bigcup_{n\in\mathbb{N}}k_{n}$, where $k_{n}$ is a finite extension of $\mathbb{F}_{p}$, $k_{n}\subseteq k_{n+1}$ and $k_{0}=\mathbb{F}_{p}$. Similarly, $\Gamma=\bigcup_{n\in\mathbb{N}}\Gamma_{n}$, where $\Gamma_{n}$ is a discrete group, $\Gamma_{n}\subseteq\Gamma_{n+1}$ and $\Gamma_{0}=\mathbb{Z}$. Let $f=[k:\mathbb{F}_{p}]$ and $e=[\Gamma:\mathbb{Z}]$; then, either $e$ or $f$ is infinite. Without loss of generality, we may assume that for each $n$, $[k_{n+1}:k_{n}][\Gamma_{n+1}:\Gamma_{n}]>1$.
@@
Let $W$ be a residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$. Then the completion $\widehat{\mathbb{Q}_{p}(X)}$ with respect to $W$ is a subfield of $\mathbb{C}_{p}$ if and only if there exists a transcendental element $\alpha$ in $\mathbb{C}_{p}$ such that $W=\mathbb{Z}_{p,\alpha}$.

### v3

@@ 2.1 Residually algebraic torsion extensions of $\mathbb{Z}_{p}$
Suppose now that $\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$.
@@
Suppose now that $n\geq 2$ and assume by induction that $(s_{n-2},s_{n-1})$ is distinguished. Let $c\in\overline{\mathbb{Q}_{p}}$ be such that $\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$. Since $(s_{n-2},s_{n-1})$ is distinguished, we have $v_{p}(s_{n-1}-c)\leq v_{p}(s_{n-1}-s_{n-2})=\delta_{n-2}<\delta_{n-1}$. Hence,
@@
We now show that $E^{\prime}=\{s_{n}\}_{n\in\mathbb{N}}$ is equivalent to $E=\{t_{n}\}_{n\in\mathbb{N}}$. Let $\{\lambda_{n}\}_{n\in\mathbb{N}}$ and $\{\delta_{n}\}_{n\in\mathbb{N}}$ be the gauges of $E$ and $E^{\prime}$, respectively. We need to show that, for each $k\in\mathbb{N}$, there exists $n\in\mathbb{N}$ such that $\lambda_{k}\leq\delta_{n}$. Since $E^{\prime}$ is unbounded, there exists $n\in\mathbb{N}$ such that $\mathbb{Q}_{p}(t_{k}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. Since $(s_{n},\delta_{n})$ is a minimal pair, we have $v_{p}(s_{n}-t_{k})<\delta_{n}$, so that
@@
Finally, we need to show that, if $k\in\mathbb{N}$, then there exist $n_{0},m_{0}\in\mathbb{N}$ such that for each $n\geq n_{0}$ and $m\geq m_{0}$, we have $v_{p}(t_{n}-s_{m})>\lambda_{k}$. Let $n_{0}$ be the smallest integer such that $\mathbb{Q}_{p}(t_{k}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n_{0}}):\mathbb{Q}_{p}]$. As in (2.6) above, $\lambda_{k}<v_{E}(X-s_{n_{0}})=\delta_{n_{0}}$. Let now $m>k$ and $n\geq n_{0}$. Then,
@@
Since $E$ is of transcendental type, for each $f\in\mathbb{Q}_{p}[X]$, $v_{E}(f)=v(f(s_{n}))$ for all $n$ sufficiently large (see §1.1). It follows that for each $\phi\in\mathbb{Q}_{p}(X)$, $\phi=f/g$, for some $f,g\in\mathbb{Q}_{p}[X]$, $v_{E}(\phi)=v_{E}(f)-v_{E}(g)$ is in $\Gamma_{n}$ for all $n$ sufficiently large. Hence, $\Gamma_{w}\subseteq\bigcup_{n}\Gamma_{n}$. Conversely, let $n\in\mathbb{N}$ and $f\in\mathbb{Q}_{p}[X]$ be of degree smaller than $\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. Then, each root $\alpha_{i}$ of $f(X)$ in $\overline{\mathbb{Q}_{p}}$ has degree smaller than $\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ and so, since $(s_{n},\delta_{n})$ is a minimal pair, we have
@@
Let now $n\in\mathbb{N}$ and $\overline{c}=\overline{f(s_{n})}\in k_{n}$, for some $f(s_{n})\in O_{n}^{*}$ where $f\in\mathbb{Q}_{p}[X]$ has degree strictly smaller than $\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. In particular, $\overline{c}\not=0$. As in (2.10), $v_{E}(f(X))=v(f(s_{m}))=0$ for each $m\geq n$. Let $\alpha_{i}$ be a root of $f(X)$ in $\overline{\mathbb{Q}_{p}}$. Then by (2.9), $\overline{v}_{E}(X-\alpha_{i})=v_{p}(s_{n}-\alpha_{i})=v_{p}(d_{i})$ for some $d_{i}\in\overline{\mathbb{Q}_{p}}$. Then
@@
Conversely, let $\phi=f/g\in W\subset\mathbb{Q}_{p}(X)$, for some $f,g\in\mathbb{Q}_{p}[X]$. Let $\alpha_{i},\beta_{j}$ be the roots in $\overline{\mathbb{Q}_{p}}$ of $f$ and $g$, respectively. There exists $n\in\mathbb{N}$ such that $\mathbb{Q}_{p}(\alpha_{i}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ and $\mathbb{Q}_{p}(\beta_{j}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ for all $i$ and $j$. Hence, as in (2.9) we have
@@
    $\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]$ and $\delta_{n}=\delta_{n}^{\prime}$.
@@
Without loss of generality, we may assume that in $\overline{\mathbb{Q}_{p}}(X)$ we have $\overline{\mathbb{Z}_{p}}_{E}=\overline{\mathbb{Z}_{p}}_{E^{\prime}}$; we let $\overline{W}=\overline{\mathbb{Z}_{p}}_{E}=\overline{\mathbb{Z}_{p}}_{E^{\prime}}$ and we denote by $w$ a valuation associated to $\overline{W}$.
We prove i). We have $s_{0},t_{0}\in\mathbb{Q}_{p}$. There exists $n\in\mathbb{N}$, $n\geq 1$, such that $w(X-s_{n-1})\leq w(X-t_{0})<w(X-s_{n})$, otherwise $t_{0}$ would be a pseudo-limit of $E$, which is not possible. In particular, $v_{p}(s_{n}-t_{0})=w(s_{n}-X+X-t_{0})=w(X-t_{0})\geq w(X-s_{n-1})=\delta_{n-1}$. If $n>1$ we have $\mathbb{Q}_{p}(t_{0}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$ so by iii) of Definition 2.1 we have that $v_{p}(s_{n}-t_{0})=v_{p}(s_{n-1}-t_{0})<v_{p}(s_{n}-s_{n-1})=\delta_{n-1}$ which is impossible. Hence, $n=1$, so $v_{p}(s_{1}-t_{0})=w(X-t_{0})\geq w(X-s_{0})$. Reversing the roles of $s_{0},t_{0}$, we get the other inequality, so $w(X-s_{0})=w(X-t_{0})=\delta_{0}=\delta_{0}^{\prime}$.
Let $n\in\mathbb{N}$ and suppose that for each $m\leq n$ we have $\mathbb{Q}_{p}(s_{m}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{m}):\mathbb{Q}_{p}]$ and $\delta_{m}=\delta_{m}^{\prime}$.
Since $\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$, by ii) of Definition 2.1 we have $v_{p}(t_{n+1}-s_{n})\leq v_{p}(t_{n+1}-t_{n})=\delta_{n}^{\prime}=\delta_{n}$. Now,
@@
Now, if $\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$, then since $(s_{n},t_{n+1})$ is distinguished, we would have $v_{p}(s_{n+1}-t_{n+1})\leq v_{p}(t_{n+1}-s_{n})$, which is impossible. Hence, $\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]\geq[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$. The other inequality is proved in a symmetrical way, so $\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n+1}):\mathbb{Q}_{p}]$.
@@
Let $n\geq 1$. If $f\in\mathbb{Q}_{p}[X]$ has degree strictly smaller than $\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]$, then by (2.10) we have $w(f(X))=v_{p}(f(s_{n}))$ and also $w(f(X))=v_{p}(f(t_{n}))$, so $v_{p}(f(s_{n}))=v_{p}(f(t_{n}))$. This proves that $\Gamma_{n}=\Gamma_{n}^{\prime}$ and so $e_{n}=e_{n}^{\prime}$.
Suppose now that $f\in\mathbb{Q}_{p}[X]$ of degree strictly smaller than $\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=[\mathbb{Q}_{p}(t_{n}):\mathbb{Q}_{p}]$ is such that $v_{p}(f(s_{n}))=v_{p}(f(t_{n}))=0$. In particular, $w(f(X))=0$ by (2.10). By (2.11) and the analogous equation where $s_{n}$ is replaced by $t_{n}$, we get that $f(s_{n}),f(t_{n})$ have the same residue as $f(X)$, so in particular, $k_{n}=k_{n}^{\prime}$. Therefore, $f_{n}=f_{n}^{\prime}$. ∎
@@ 2.2 Residually algebraic extensions of $\mathbb{Z}_{p}$ which are DVRs
In this section we characterize DVRs of $\mathbb{Q}_{p}(X)$ extending $\mathbb{Z}_{p}$ such that the residue field extension is algebraic, necessarily of infinite degree by Corollary 2.12; this fact has already been noted in a different way in [19, p. 4217]. We will see in §2.3 that there is no such restriction on the residue field degree for DVRs of $\mathbb{Q}(X)$ which are residually algebraic extensions of $\mathbb{Z}_{(p)}$ (see Corollary 2.28).
@@
[…]
@@
[…]
@@
The following result is analogous to [19, Theorem 2.5].
@@
We show in Theorem 2.21 that given any algebraic extension $k$ of $\mathbb{F}_{p}$ and group $\Gamma$ such that $\mathbb{Z}\subseteq\Gamma\subseteq\mathbb{Q}$, there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $\mathbb{Z}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$, provided that either $k:\mathbb{F}_{p}]$ is infinite or $\Gamma$ is not discrete (this condition being necessary by Corollary 2.12).
@@
Let $k$ be an algebraic extension of $\mathbb{F}_{p}$ and $\Gamma$ a totally ordered group with $\mathbb{Z}\subseteq\Gamma\subseteq\mathbb{Q}$, such that either $k:\mathbb{F}_{p}]$ or $\Gamma:\mathbb{Z}]$ is infinite (the last condition is equivalent to $\Gamma$ being not discrete). Then there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $k_{\alpha}=k$ and $\Gamma_{\alpha}=\Gamma$. In particular, $\mathbb{Z}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$.
Note that, by Corollary 2.12, the last claim shows that $k:\mathbb{F}_{p}]\cdot[\Gamma:\mathbb{Z}]=\infty$ is necessary.
@@
Since $\overline{\mathbb{F}_{p}}$ is countable, we may suppose that $k=\bigcup_{n\in\mathbb{N}}k_{n}$, where $k_{n}$ is a finite extension of $\mathbb{F}_{p}$, $k_{n}\subseteq k_{n+1}$ and $k_{0}=\mathbb{F}_{p}$. Similarly, $\Gamma=\bigcup_{n\in\mathbb{N}}\Gamma_{n}$, where $\Gamma_{n}$ is a discrete group, $\Gamma_{n}\subseteq\Gamma_{n+1}$ and $\Gamma_{0}=\mathbb{Z}$. Let $f=[k:\mathbb{F}_{p}]$ and $e=[\Gamma:\mathbb{Z}]$; then, either $e$ or $f$ is infinite. Without loss of generality, we may assume that for each $n$, $k_{n+1}:k_{n}][\Gamma_{n+1}:\Gamma_{n}]>1$.
@@
Let $W$ be a residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$. Then the completion $\widehat{\mathbb{Q}_{p}(X)}$ with respect to $W$ is (isomorphic to) a subfield of $\mathbb{C}_{p}$ if and only if there exists a transcendental element $\alpha$ in $\mathbb{C}_{p}$ such that $W=\mathbb{Z}_{p,\alpha}$.